# Simulating stochastic population dynamics: The Linear Noise Approximation can capture non-linear phenomena

### v3

@@ 6 Discussion
Let $T^{c}$ be the eigenspace corresponding to the $n_{0}+1$ eigenvalues of (27) with zero real part. Then, there exists a smooth $(n_{0}+1)$-dimensional manifold $\mathcal{W}^{c}$ defined locally to ${\bm{x}}={\bm{x}}_{\alpha_{p}}$ and $\alpha=\alpha_{p}$ and tangent to $T^{c}$ at ${\bm{x}}={\bm{x}}_{\alpha_{p}}$ and $\alpha=\alpha_{p}$.
Moreover since $\dot{\alpha}=0$ the hyperplanes, $\Pi_{\alpha^{\prime}}=\{({\bm{x}},\alpha)\ |\ \alpha=\alpha^{\prime}\}$, $\alpha^{\prime}\in\mathbb{R}$ are invariant with respect to (26) and so, for each $\alpha$ local to $\alpha_{p}$, $\mathcal{W}^{c}$ is foliated by the $n_{0}$ dimensional manifolds $\mathcal{W}^{c}_{\alpha}=\mathcal{W}^{c}\cap\Pi_{\alpha}.\$For each $\alpha$ local to $\alpha_{p}$ we call $\mathcal{W}^{c}_{\alpha}$ the centre manifold.
Finally, for each $\alpha$ local to $\alpha_{p}$, there exists a neighbourhood $\Delta({\bm{x}}_{\alpha_{p}})$ of ${\bm{x}}={\bm{x}}_{\alpha_{p}}$ such that if a path of (6) starting in $\Delta({\bm{x}}_{\alpha_{p}})$, remains in $\Delta({\bm{x}}_{\alpha_{p}})$, then this path approach $\mathcal{W}^{c}_{\alpha}$ as $t\to\infty$.
In other words, the Center Manifold Theorem states that there exists a parameter-dependent $n_{0}$-dimensional invariant center manifold $\mathcal{W}^{c}_{\alpha}$ which is locally attracting. The next theorem (see [52]) describes a decomposition of non-hyperbolic systems into a linear and a non-linear component.
Theorem 2 (Shoshitaishvili Reduction Principle).
Let $\alpha$ be local to $\alpha_{p}$ and let $({\bm{u}},{\bm{v}},\bm{w})$ be a coordinate system with coordinates, ${\bm{u}}\in\mathbb{R}^{n_{0}},\ {\bm{v}}\in\mathbb{R}^{n_{-}}$ and $\bm{w}\in\mathbb{R}^{n_{+}}$ lying on the eigenspaces corresponding to the $_{0}, n+n_{+}$, $n_{-}$ eigenvalues of $\bm{J}({\bm{x}}_{\alpha},\alpha)$ with zero, positive and negative real part, when $\alpha=\alpha_{p}$, respectively.
Then, the centre manifold $\mathcal{W}^{c}_{\alpha}$ can be locally represented as a graph of a smooth function ${\bm{V}}_{\alpha}:\mathbb{R}^{n_{0}}\to\mathbb{R}^{n_{-}+n_{+}}$ with ${\bm{V}}_{\alpha}({\bm{u}})=O(\|{\bm{u}}\|^{2})$ [^4] This implies that the Taylor expansion of ${\bm{V}}_{\alpha}({\bm{u}})$ starts with at least quadratic terms.;
$\mathcal{W}^{c}_{\alpha}=\{({\bm{u}},{\bm{v}},\bm{w})\ |\ ({\bm{v}},\bm{w})={\bm{V}}_{\alpha}({\bm{u}})\}.$
Moreover, near the equilibrium ${\bm{x}}_{p}$, (6) is locally topologically equivalent to the following system;
$\left\{\begin{aligned} \dot{{\bm{u}}}&={\bm{A}}{\bm{u}}+{\bm{g}}({\bm{u}},{\bm{V}}_{\alpha}({\bm{u}})),\\ \dot{{\bm{v}}}&=-{\bm{v}},\\ \dot{\bm{w}}&=\bm{w},\end{aligned}\right$. (28)
near its equilibrium. Here $\bm{A}$} is an $n_{0}\times n_{0}$ matrix with all its eigenvalues, for $\alpha=\alpha_{p}$, on the imaginary axis and ${\bm{g}}:\mathbb{R}^{n}\to\mathbb{R}^{n_{0}}$ is a smooth function with Taylor expansions starting with at least quadratic terms.
The first equation in (28) is a restriction of equation (26) on the center manifold $\mathcal{W}^{c}_{\alpha}$. The other two equations in (28) present linear, transient dynamics. Therefore, for small $|\alpha-\alpha_{p}|$, all non-transient events are captured on the invariant $\mathcal{W}^{c}_{\alpha}$ by the $n_{0}$-dimensional system of $\bm{u}$}, where $n_{0}$ can be much smaller than $n$. It is worth noting that if $n_{+}=0$, the third equation vanishes, and the center manifold $\mathcal{W}^{c}_{\alpha}$ is attracting.
@@
$k_{t2a}$ degradation of IkBa (IKK dependent from trimer) 0.1 $s^{-1}$
@@
The reaction rates used for the SSA are provided in the table below. The values of the parameters are the same as in Table S2 with suitable rate constant conversions applied.
@@
See Figures S9, S10, S11.
@@
If, for each $t\geq 0$, we let ${\bm{Y}}(t)=(Y_{1}(t),Y_{2}(t))^{T}$ denote the number of molecules of $A$ and $B$ respectively, then these reactions occur according to the following intensities
@@
From studying the Jacobian matrix of (29), we may conclude the system has a non-hyperbolic equilibrium at $(x_{1},x_{2})=(1,2)^{T}$, when $b=2$ and $c=1$, with purely imaginary conjugate eigenvalues $\pm i$. Through small deviations of $b$ we deduce the Hopf bifurcation is supercritical with losses of stability occurring as the eigenvalues cross the imaginary axis. Parameters $b<2$ give rise to a stable equilibrium, whereas $b>2$ gives rise to an unstable equilibrium surrounded by a unique and stable limit cycle.
@@
Table S8: Reactions and their corresponding propensity functions for SSA simulations.
@@
Figure S12: Comparison between SSA (red) and pcLNA (blue) for the Brusselator system at $\Omega=1000$ with parameter $b=1.7$. Panels (A)–(D) correspond to times $t=\tau+0.5,\ 3\tau+0.5,\ 5\tau+0.5,\ 7\tau+0.5$ with period $\tau=6.2919$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are given in Tables S5 and S6.
Figure S13: Comparison between SSA (red) and pcLNA (blue) for the Brusselator system at $\Omega=1000$ with parameter $b=2$. Panels (A)–(D) correspond to times $t=0.5,\ 2\tau+0.5,\ 4\tau+0.5,\ 6\tau+0.5$ with period $\tau=6.35$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are provided in Tables S5 and S6.
Figure S14: Comparison between SSA (red) and pcLNA (blue) for the Brusselator system at $\Omega=1000$ with parameter $b=2.3$. Panels (A)–(D) correspond to times $t=\tau+0.5,\ 3\tau+0.5,\ 5\tau+0.5,\ 7\tau+0.5$ with period $\tau=6.4276$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are listed in Tables S5 and S6.

### v4

@@ 7 Discussion
Let $E^{c}$ denote the eigenspace of $\mathcal{A}$ corresponding to the $n_{0}+1$ eigenvalues with zero real part. Then there exists a smooth $(n_{0}+1)$-dimensional invariant manifold $\mathcal{W}^{c}$, defined locally near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$ and tangent to $E^{c}$ at that point.
Since $\dot{\alpha}=0$, the hyperplanes
$\Pi_{\alpha^{\prime}}:=\{({\bm{x}},\alpha)\mid\alpha=\alpha^{\prime}\},\qquad\alpha^{\prime}\in\mathbb{R},$
are invariant under (40). Consequently, for each $\alpha$ near $\alpha_{p}$, the center manifold is foliated by $n_{0}$-dimensional invariant manifolds
$\mathcal{W}^{c}_{\alpha}:=\mathcal{W}^{c}\cap\Pi_{\alpha},$
which we call the centre manifolds. Moreover, for each such $\alpha$, there exists a neighbourhood $\Delta({\bm{x}}_{\alpha_{p}})$ such that any trajectory of (40) starting in $\Delta({\bm{x}}_{\alpha_{p}})$ either converges to $\mathcal{W}^{c}_{\alpha}$ or leaves $\Delta({\bm{x}}_{\alpha_{p}})$ exponentially fast as $t\to\infty$.
Takens normal form for the extended system
Let $\lambda_{1},\dots,\lambda_{n_{cu}}$, where $n_{cu}=n_{0}+n_{-}+1$, denote the eigenvalues of $\mathcal{A}$ with zero or negative real part including the zero eigenvalue associated with $\dot{\alpha}=0$. For each integer $k\geq 1$, we say that (40) satisfies the $k$-Sternberg nonresonance condition (see [62, Definition 2 and Section 4]) if the following holds: for any choice of integers $c_{1},\dots,c_{n_{cu}}$ satisfying
$2\leq\sum_{i=1}^{n_{cu}}c_{i}\leq\gamma(\mathcal{A},k),$
the sums,
$\sum_{i=1}^{n_{cu}}c_{i}\,\lambda_{i}\neq 0,\qquad-\lambda_{j}+\sum_{i=1}^{n_{cu}}c_{i}\,\lambda_{i}\neq 0\quad\text{for all }j=1,\dots,n_{cu}.$
Here the integer $\gamma(\mathcal{A},k)$ is defined as follows. Here we use the notation $\Re(\mu)$ to denote the real part of an eigenvalue $\mu$. Let $\mu_{1},\dots,\mu_{n_{-}+n_{+}}$ denote the eigenvalues of $\mathcal{A}$ with non zero real part, ordered so that
$\Re(\mu_{1})\leq\cdots\leq\Re(\mu_{n_{-}})<0<\Re(\mu_{n_{-}+1})\leq\cdots\leq\Re(\mu_{n_{-}+n_{+}}).$
Define
$\overline{M}:=\max_{i>n_{-}}\Re(\mu_{i}),\qquad\overline{m}:=\min_{i>n_{-}}\Re(\mu_{i}),\qquad\overline{N}:=\max_{i\leq n_{-}}|\Re(\mu_{i})|,\qquad\overline{n}:=\min_{i\leq n_{-}}|\Re(\mu_{i})|.$
Then $\beta(\mathcal{A},k)$ is the smallest integer such that
$\overline{N}+r\,\overline{M}-(r-\beta(\mathcal{A},k))\,\overline{n}<0\qquad\text{for all integers }r\leq k,$
and $\gamma(\mathcal{A},k)$ is the smallest integer such that
$\overline{M}+r\,\overline{N}-(r-\gamma(\mathcal{A},k))\,\overline{m}<0\qquad\text{for all integers }r\leq\beta(\mathcal{A},k).$
The above conditions hold for almost all values of parameter $\alpha$. In reaction networks, the free parameters $\alpha$ are typically reaction rate constants, for which the eigenvalues $\lambda_{i}$ of the Jacobian depend smoothly. Resonance requires exact relations of the form $\sum_{i}c_{i}\lambda_{i}=0$ for integers $c_{i}$, which occur only on sets of Lebesgue measure zero in parameter space and are destroyed by arbitrarily small parameter perturbations. Consequently, for analytical purposes the nonresonance conditions are satisfied almost surely, and for practical simulation purposes they can be enforced by arbitrarily small perturbations of the parameter $\alpha$ that do not affect the qualitative dynamics.
The following theorem requires these conditions to hold for all integers $k\geq 2$. It is therefore sufficient—and in practice straightforward—to verify that they hold for a given system for the cases $k=1,2$.
For the remaninder of this section, assume that (40) satisfies the $k$-Sternberg nonresonance condition for some integer $k\geq 2$. Then the following theorem, first proved in [62, Section 4], holds. Below we use the notation $C^{m}$, $m\in\{1,2,\dots\}$, to denote the set of functions that have a $m$-th derivative that is continuous in its domain.
Theorem 2 (Takens).
Suppose $\bm{F}$} in (40) is $C^{\infty}$ and the above $k$-Sternberg nonresonance condition holds for some $k\geq 2$. Then there exists a local $C^{k}$ change of coordinates
$({\bm{x}},\alpha)\mapsto({\bm{u}},{\bm{v}},\bm{w},\alpha),\qquad{\bm{u}}\in\mathbb{R}^{n_{0}},\ {\bm{v}}\in\mathbb{R}^{n_{-}},\ \bm{w}\in\mathbb{R}^{n_{+}},$
defined near $({\bm{x}}_{\alpha_{p}},\alpha_{p})$, such that the centre manifold $\mathcal{W}^{c}_{\alpha}$ in Theorem 1 is mapped precisely onto the linear subspace
$\{({\bm{u}},{\bm{v}},\bm{w},\alpha):\ {\bm{v}}=\bm{w}=0\},$
and the extended system (40) takes the standard form
$\left\{\begin{aligned} \dot{\bm{u}}&=\bm{f}({\bm{u}},\alpha),\\ \dot{\bm{v}}&={\bm{A}}_{\text{s}}({\bm{u}},\alpha){\bm{v}},\\ \dot{\bm{w}}&={\bm{A}}_{\text{u}}({\bm{u}},\alpha)\bm{w},\\ \dot{\alpha}&=0,\end{aligned}\right$. (42)
where $\bm{f},{\bm{A}}_{\text{s}},{\bm{A}}_{\text{u}}$ are $C^{k}$ functions, all eigenvalues of the $n_{-}\times n_{-}$ matrix ${\bm{A}}_{\text{s}}({\bm{u}}(t),\alpha)$ have negative real part, and all eigenvalues of the $n_{+}\times n_{+}$ matrix ${\bm{A}}_{\text{u}}({\bm{u}}(t),\alpha)$ have positive real part.
In particular, the reduced dynamics on $\mathcal{W}^{c}_{\alpha}$ is represented solely by the evolution of the centre coordinates $\bm{u}$}, while the stable and unstable coordinates $({\bm{v}},\bm{w})$ evolve according to the linear equations determined by ${\bm{A}}_{\text{s}}({\bm{u}},\alpha)$ and ${\bm{A}}_{\text{u}}({\bm{u}},\alpha)$.
Relation to local topological equivalence
From Theorem 2, there exist neighbourhoods
$\mathcal{E}\subset\mathbb{R}^{n}\quad\text{of }{\bm{x}}_{\alpha_{p}},\qquad\mathcal{F}\subset\mathbb{R}^{n_{0}+n_{-}+n_{+}}\quad\text{of the origin},$
and an interval $\mathcal{I}\subset\mathbb{R}$ containing $\alpha_{p}$, such that the coordinate transformation
$({\bm{x}},\alpha)\longmapsto({\bm{u}},{\bm{v}},\bm{w},\alpha)$
is a $C^{k}$ diffeomorphism from $\mathcal{E}\times\mathcal{I}$ onto $\mathcal{F}\times\mathcal{I}$.
For each fixed $\alpha\in\mathcal{I}$, restricting this transformation to the invariant hyperplane
$\Pi_{\alpha}=\{({\bm{x}},\alpha)\in\mathbb{R}^{n+1}\}$
in Theorem 1, defines a $C^{k}$ diffeomorphism
${\bm{h}}_{\alpha}:\ \mathcal{E}_{\alpha}\longrightarrow\mathcal{F}_{\alpha},\qquad{\bm{x}}\longmapsto({\bm{u}},{\bm{v}},\bm{w}),$
where
$\mathcal{E}_{\alpha}:=\mathcal{E}\cap\{{\bm{x}}:\ ({\bm{x}},\alpha)\in\Pi_{\alpha}\},\quad\text{and}\quad\mathcal{F}_{\alpha}:={\bm{h}}_{\alpha}(\mathcal{E}_{\alpha}).$
Under this mapping, trajectories of the original system (6) in $\mathcal{E}_{\alpha}$ correspond to trajectories of the standard-form system (42) in $\mathcal{F}_{\alpha}$ with the same time orientation. Consequently, for each $\alpha\in\mathcal{I}$, the map ${\bm{h}}_{\alpha}$ realises the local topological equivalence of Definition 1 between the original system (6) in $\mathcal{E}_{\alpha}$ and the standard-form system (42) in $\mathcal{F}_{\alpha}$.
Eigenspace projection approximation of the centre coordinates
Although ${\bm{h}}_{\alpha}$ mapping the original coordinates $\bm{x}$} to the centre–stable–unstable coordinates $({\bm{u}},{\bm{v}},\bm{w})$ is in general nonlinear, its action on the centre coordinates admits a simple first–order approximation near the non–hyperbolic equilibrium ${\bm{x}}_{\alpha_{p}}$. By construction, the centre manifold $\mathcal{W}^{c}_{\alpha}$ is tangent at ${\bm{x}}_{\alpha_{p}}$ to the eigenspace $E^{c}$ associated with the eigenvalues of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ having zero real part. Consequently, the differential $\nabla_{x}h_{\alpha}({\bm{x}}_{\alpha_{p}})$ restricted to the centre coordinates coincides with the linear projection onto $E^{c}$.
Let ${\bm{R}}_{c}\in\mathbb{R}^{n\times n_{0}}$ denote a matrix whose columns form a basis of this eigenspace, and define the associated projection $\bm{P}_{c}:={\bm{R}}_{c}^{-1}$ onto the centre coordinates. Then, for any state $\bm{X}$} sufficiently close to ${\bm{x}}_{\alpha_{p}}$, the centre component $\bm{U}$} of $h_{\alpha}({\bm{X}})$ satisfies
${\bm{U}}=P_{c}({\bm{X}}-{\bm{x}}_{\alpha_{p}})+O\left(\|{\bm{X}}-{\bm{x}}_{\alpha_{p}}\|^{2}\right).$ (43)
This approximation shows that, to leading order, the nonlinear centre coordinates are obtained simply by projecting deviations from the equilibrium onto $E^{c}$.
Application of Itô’s Lemma to the image of the LNA under the local homeomorphism
To simplify notation as before, in this section we use the notation $x_{t}=x(t)$, for all states considered (i.e. $\bm{x}$}, $\bm{\xi}$, $\bm{X}$}). Under the non-resonance conditions on the eigenvalues of $\bm{J}({\bm{x}}_{\alpha_{p}},\alpha_{p})$ required for Theorem 2, the mapping ${\bm{h}}_{\alpha}:\mathcal{E}_{\alpha}\to\mathcal{F}_{\alpha}$ is $C^{k}$ for some $k\geq 2$. Given that ${\bm{X}}_{0}={\bm{x}}_{0}+\Omega^{-1/2}\bm{\xi}_{0}\in\mathcal{E}_{\alpha}$, this level of regularity is sufficient to apply Itô’s Lemma (see [38, Theorem 5.10]) to the image of the continuous semimartingale ${\bm{X}}_{t}={\bm{x}}_{t}+\Omega^{-1/2}\bm{\xi}_{t}$ under ${\bm{h}}_{\alpha}$ and obtain an SDE representation of its evolution for times $t\in[0,\tau)$ where,
$\tau:=\inf\{t\geq 0:{\bm{X}}_{t}={\bm{x}}_{t}+\Omega^{-1/2}\bm{\xi}_{t}\notin\mathcal{E}_{\alpha}\}.$
To this end,
$d{\bm{h}}_{\alpha}({\bm{X}}_{t})$ $=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{X}}_{t})\ d{\bm{X}}_{t}+\frac{1}{2\Omega}\text{Tr}\left({\bm{E}}_{t}^{\top}\nabla^{2}_{\bm{x}}{\bm{h}}_{\alpha}({\bm{X}}_{t}){\bm{E}}_{t}\right)\ dt$ (44)
$=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{X}}_{t})\left({\bm{F}}({\bm{x}}_{t},\alpha)+\frac{1}{\sqrt{\Omega}}\bm{J}_{t}\bm{\xi}_{t}\right)\ dt$
$\qquad+\frac{1}{\sqrt{\Omega}}\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{X}}_{t}){\bm{E}}_{t}\ d{\bm{B}}_{t}+\frac{1}{2\Omega}\text{Tr}\left({\bm{E}}_{t}^{\top}\nabla^{2}_{\bm{x}}{\bm{h}}_{\alpha}({\bm{X}}_{t}){\bm{E}}_{t}\right)\ dt.$
Taylor expanding $\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{X}}_{t})=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}+\Omega^{-1/2}\bm{\xi}_{t})$ about ${\bm{x}}_{t}$ gives
$\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{X}}_{t})=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t})+\frac{1}{\sqrt{\Omega}}\bm{\xi}_{t}^{\top}\nabla^{2}_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t})+O(\Omega^{-1}),$
and so (44) becomes,
$d{\bm{h}}_{\alpha}({\bm{X}}_{t})$ $=\Bigg(\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}){\bm{F}}({\bm{x}}_{t},\alpha)+\frac{1}{\sqrt{\Omega}}\bm{\xi}^{\top}_{t}\nabla^{2}_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}){\bm{F}}({\bm{x}}_{t},\alpha)$ (45)
$\qquad+\frac{1}{\sqrt{\Omega}}\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t})\bm{J}_{t}\bm{\xi}_{t}\Bigg)\ dt+\frac{1}{\sqrt{\Omega}}\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}){\bm{E}}_{t}\ d{\bm{B}}_{t}+O(\Omega^{-1}).$
By Theorem 2, $({\bm{u}}_{t},{\bm{v}}_{t},\bm{w}_{t})^{\top}:={\bm{h}}_{\alpha}({\bm{x}}_{t})$ evolves according to (42) and hence
$\frac{d}{dt}({\bm{u}}_{t},{\bm{v}}_{t},\bm{w}_{t})=\frac{d}{dt}{\bm{h}}_{\alpha}({\bm{x}}_{t})=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t})\frac{d}{dt}{\bm{x}}_{t}=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}){\bm{F}}({\bm{x}}_{t},\alpha),$ (46)
which enables to write the first term in the RHS of (45)
$\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}){\bm{F}}({\bm{x}}_{t},\alpha)=\frac{d}{dt}{\bm{h}}_{\alpha}({\bm{x}}_{t})=\begin{pmatrix}\bm{f}({\bm{u}}_{t},\alpha)\\ {\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha){\bm{v}}_{t}\\ {\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{w}_{t}\end{pmatrix}=\tilde{\bm{F}}({\bm{h}}_{\alpha}({\bm{x}}_{t}))$ (47)
where
$\tilde{\bm{F}}:\ \begin{pmatrix}{\bm{u}}\\ {\bm{v}}\\ \bm{w}\end{pmatrix}\longmapsto\begin{pmatrix}\bm{f}({\bm{u}},\alpha)\\ {\bm{A}}_{\text{s}}({\bm{u}},\alpha){\bm{v}}\\ {\bm{A}}_{\text{u}}({\bm{u}},\alpha)\bm{w}\end{pmatrix}$. (48)
Then, using the product rule of differentiation, the second and third term in (45) are
$\bm{\xi}^{\top}_{t}\nabla^{2}_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}){\bm{F}}({\bm{x}}_{t},\alpha)$ $+\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t})\bm{J}_{t}\bm{\xi}_{t}$
$\qquad=\nabla_{\bm{x}}\left(\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}){\bm{F}}({\bm{x}}_{t},\alpha)\right)\bm{\xi}_{t}$
$\qquad=\nabla_{({\bm{u}},{\bm{v}},\bm{w})}\left(\widehat{\bm{F}}({\bm{h}}_{\alpha}({\bm{x}}_{t})\right)\bm{\xi}_{t}$
$\qquad=\begin{pmatrix}\nabla_{\bm{u}}\bm{f}({\bm{u}}_{t},\alpha)&0&0\\ \nabla_{\bm{u}}({\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha){\bm{v}}_{t})&{\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha)&0\\ \nabla_{\bm{u}}({\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{w}_{t})&0&{\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\end{pmatrix}\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t})\bm{\xi}_{t}$
$\qquad=\begin{pmatrix}\nabla_{\bm{u}}\bm{f}({\bm{u}}_{t},\alpha)\bm{\xi}_{t}^{{\bm{u}}}\\ \nabla_{\bm{u}}({\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha){\bm{v}}_{t})\bm{\xi}_{t}^{{\bm{u}}}+{\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha)\bm{\xi}_{t}^{{\bm{v}}}\\ \nabla_{\bm{u}}({\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{w}_{t})\bm{\xi}_{t}^{{\bm{u}}}+{\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{\xi}_{t}^{\bm{w}}\end{pmatrix}$ (49)
where $(\bm{\xi}^{{\bm{u}}}_{t},\bm{\xi}^{{\bm{v}}}_{t},\bm{\xi}^{\bm{w}}_{t})^{\top}:=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t})\bm{\xi}_{t}$. Finally, by substituting (47) and (S3) into (45) we have,
$d{\bm{h}}_{\alpha}({\bm{X}}_{t})=\begin{pmatrix}\bm{f}({\bm{u}}_{t},\alpha)\\ {\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha){\bm{v}}_{t}\\ {\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{w}_{t}\end{pmatrix}\ dt+\frac{1}{\sqrt{\Omega}}\begin{pmatrix}\nabla_{\bm{u}}\bm{f}({\bm{u}}_{t},\alpha)\bm{\xi}_{t}^{{\bm{u}}}\\ \nabla_{\bm{u}}({\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha){\bm{v}}_{t})\bm{\xi}_{t}^{{\bm{u}}}+{\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha)\bm{\xi}_{t}^{{\bm{v}}}\\ \nabla_{\bm{u}}({\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{w}_{t})\bm{\xi}_{t}^{{\bm{u}}}+{\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{\xi}_{t}^{\bm{w}}\end{pmatrix}\ dt\\ +\frac{1}{\sqrt{\Omega}}\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{X}}_{t}){\bm{E}}_{t}\ d{\bm{B}}_{t}+O(\Omega^{-1}).$ (50)
From inspecting the form of (50), we conclude that
${\bm{h}}_{\alpha}({\bm{X}}_{t})=\begin{pmatrix}{\bm{u}}_{t}\\ {\bm{v}}_{t}\\ \bm{w}_{t}\end{pmatrix}+\frac{1}{\sqrt{\Omega}}\begin{pmatrix}\bm{\xi}_{t}^{{\bm{u}}}\\ \bm{\xi}_{t}^{{\bm{v}}}\\ \bm{\xi}_{t}^{\bm{w}}\end{pmatrix}+O(\Omega^{-1}),$ (51)
where ${\bm{u}}_{t},{\bm{v}}_{t},\bm{w}_{t}$ evolve according to (42) and $\bm{\xi}^{{\bm{u}}}_{t},\bm{\xi}^{{\bm{v}}}_{t},\bm{\xi}^{\bm{w}}_{t}$ evolve according to the following system of linear SDEs
$\left\{\begin{aligned} d\bm{\xi}^{{\bm{u}}}_{t}&=\nabla_{\bm{u}}\bm{f}({\bm{u}}_{t},\alpha)\bm{\xi}^{{\bm{u}}}_{t}\ dt+{\bm{H}}^{\bm{u}}_{t}\ d{\bm{B}}_{t},\\ d\bm{\xi}^{{\bm{v}}}_{t}&=\left(\nabla_{\bm{u}}({\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha){\bm{v}}_{t})\bm{\xi}^{{\bm{u}}}_{t}+{\bm{A}}_{\text{s}}({\bm{u}}_{t},\alpha)\bm{\xi}^{{\bm{v}}}_{t}\right)\ dt+{\bm{H}}^{\bm{v}}_{t}\ d{\bm{B}}_{t},\\ d\bm{\xi}^{\bm{w}}_{t}&=\left(\nabla_{\bm{u}}({\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{w}_{t})\bm{\xi}^{\bm{u}}_{t}+{\bm{A}}_{\text{u}}({\bm{u}}_{t},\alpha)\bm{\xi}^{\bm{w}}_{t}\right)\ dt+{\bm{H}}^{\bm{w}}_{t}\ d{\bm{B}}_{t},\end{aligned}\right$. (52)
with
$\begin{bmatrix}{\bm{H}}^{\bm{u}}_{t}\\ {\bm{H}}^{\bm{v}}_{t}\\ {\bm{H}}^{\bm{w}}_{t}\end{bmatrix}:=\nabla_{\bm{x}}{\bm{h}}_{\alpha}({\bm{x}}_{t}){\bm{E}}_{t}.$ (53)
@@
In this section, we give further details related to the numerical investigations included in Section 6. In particular, we provide the details of the systems used and additional comparisons between SSA and pcLNA.
@@
$k_{t2a}$ degradation of IkBa (IKK dependent) 0.1 $s^{-1}$
@@
The propensity functions used for the SSA are provided in Table S5. The values of the parameters are the same as in Table S3 with suitable rate constant conversions applied.
@@
See caption of Figures S9, S10, S11 for more details on additional comparisons for this system, in the cases where the parameter values give a system that presents a stable focus (see Figure S9), the case where the parameter values are exactly on the Hopf bifurcation point (see Figure S10), and the case where a limit cycle appears (see Figure S11).
@@
If, for each $t\geq 0$, we let ${\bm{Y}}(t)=(Y_{1}(t),Y_{2}(t))^{\top}$ denote the number of molecules of $A$ and $B$ respectively, then these reactions occur according to the following intensities
@@
From studying the Jacobian matrix of (54), we may conclude the system has a non-hyperbolic equilibrium at $(x_{1},x_{2})=(1,2)^{\top}$, when $b=2$ and $c=1$, with purely imaginary conjugate eigenvalues $\pm i$. Through small deviations of $b$ we deduce the Hopf bifurcation is supercritical with losses of stability occurring as the eigenvalues cross the imaginary axis. Parameters $b<2$ give rise to a stable equilibrium, whereas $b>2$ gives rise to an unstable equilibrium surrounded by a unique and stable limit cycle.
@@
Table S9: Brusselator reactions and their corresponding propensity functions for SSA simulations.
@@
Figure S12: Comparison between SSA (red) and pcLNA (blue) for the Brusselator system at $\Omega=1000$ with parameter $b=1.7$. Panels (A)–(D) correspond to times $t=\tau+0.5,\ 3\tau+0.5,\ 5\tau+0.5,\ 7\tau+0.5$ with period $\tau=6.2919$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are given in Tables S6 and S7.
Figure S13: Comparison between SSA (red) and pcLNA (blue) for the Brusselator system at $\Omega=1000$ with parameter $b=2$. Panels (A)–(D) correspond to times $t=\tau+0.5,\ 3\tau+0.5,\ 5\tau+0.5,\ 7\tau+0.5$ with period $\tau=6.35$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are provided in Tables S6 and S7.
Figure S14: Comparison between SSA (red) and pcLNA (blue) for the Brusselator system at $\Omega=1000$ with parameter $b=2.3$. Panels (A)–(D) correspond to times $t=\tau+0.5,\ 3\tau+0.5,\ 5\tau+0.5,\ 7\tau+0.5$ with period $\tau=6.4276$. Each panel shows (left) the states recorded at time $t$ of the simulations with solutions of the corresponding RRE, and (right) the corresponding empirical probability density functions of each system variable. Initial conditions and parameter values are listed in Tables S6 and S7.